OpenStudy (anonymous):

MEDAL A lake was measured with an optical laser ranging instrument as shown in the drawing. Its scale is 1 cm = 1 km. How wide is the lake? 46.5 m 465 m 4,650 m 46,500 m

OpenStudy (anonymous):



OpenStudy (anonymous):

@Mr.ClayLordMath

OpenStudy (anonymous):

@dpasingh @Hero

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):



OpenStudy (anonymous):

Let BC be the width of the lake so \[\tan 30^0 = \frac {AB}{BC} \rightarrow \frac{1}{\sqrt{3}} = \frac {4.65}{BC}\] \[\frac{1}{\sqrt{3}} = \frac {4.65}{BC} \] \[BC=4.65\sqrt{3} =4.65 \times 1.73 =8.0445 \approx 8\] Hence width of the lake is 8 KM

OpenStudy (anonymous):

Your ans is not matching with waht i calculated, pls chek the figure properly and then report me.

OpenStudy (anonymous):

mmmm.i think that the answer don't need to have the same value meters have : my answer was option 3

OpenStudy (anonymous):

just see the scale 1cm = 1 km

OpenStudy (anonymous):

so 1 cm = 1000m

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so!!!!

OpenStudy (anonymous):

According to given info the width of the lake must be 8 KM

OpenStudy (anonymous):

OMG!! we dont have this answer

OpenStudy (anonymous):

I am pretty sure the answer would be 4,650 because since 1cm=1km, you would convert 4,65cm into 4.65km. And 4.65km=4,650m :) hope I helped